OpenStudy (anonymous):

[2.06] Solve the equation 1 over 4 x + 2 = 5 for x. 28 20 14 12

OpenStudy (anonymous):

can you write the equation using parentheses please.

OpenStudy (anonymous):

I can help but I'm not exactly sure what the equation is.

OpenStudy (anonymous):

ya its kind of hard

OpenStudy (anonymous):

1/4x+2=5 subtract both of the sides by two -2 -2 1/4x=5-2 1/4x=3 divide both sides by 1/4 . /1/4 /1/4 x = 3 / 1/4 3 divided by .25 is 12 x = 12

OpenStudy (anonymous):

It's not....I just don't know what's the exact equation. only you do.

OpenStudy (anonymous):

1/(4x+2)=5 can also be an equation....with a different answer.

OpenStudy (anonymous):

that is if x isn't in denominator

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

I just think this guy has no idea what he is doing lol

OpenStudy (anonymous):

thats so funny noooooooooo

OpenStudy (wolf1728):

I can see what RyanL is talking about. Is the attachment the way the equation should read?

OpenStudy (wolf1728):

Or is this the way the equation should read?

OpenStudy (anonymous):

@littlelotte:end the answering ???????? @wolf1728:???

OpenStudy (wolf1728):

Hello Eali

OpenStudy (anonymous):

hi friend ... Hello to all ..:) The @supanick14 got it?:)

OpenStudy (anonymous):